OpenStudy (anonymous):

Jose is paying for a meal using bills in his wallet. He has five 1 bills and four 5 bills. If he selects two bills at random, one at a time from his wallet, without replacement, what is the probability that he will choose two 5 bills to pay for the meal? Show your work.

OpenStudy (anonymous):

4/9*3/8=12/72=1/6